# Hypothetical Lorentz invariance violation and the muon content of extensive air showers

### v1

@@ I Introduction
One of the most intriguing problems in cosmic-ray physics is that of the “muon excess” in extensive atmospheric showers (EAS), or “the muon puzzle”. According to data from a number of experiments, the muon content in EAS exceeds significantly the predictions of models of the development of the shower for the same energy of the primary particle. Solving the muon excess problem is a key to determining accurate models of the development of EAS. Understanding the EAS development is a necessary prerequisite for any reliable study of ultrahigh energy cosmic rays, and at the same time, EAS physics is a tool to study particle interactions in kinematical regimes hard to probe at accelerators.
@@ II.1 An EFT approach to $n=2$ LIV Quantum Electrodynamics
Thus, analyzing LIV for mass scales $10^{(13\dots 16)}$ GeV we can neglect LIV for electrons. In this paper, we neglect LIV for electrons even for higher mass scales since we propose a particular LIV solution of muon puzzle which can be later generalized to account for electron LIV parameters.
@@ IV Monte-Carlo simulations
One can see that the Monte–Carlo simulations explicitly justify the use of Eqs. (11)–(14), and (17). The actual particle ratios trend is well captured by our parametric model.
Figure 2: LI–to–LIV ratios for the ground-level average number of particles, test data set (main figures: electrons, inset figures: muons). The colored points and the error bars correspond to MC simulations (proton-induced EASs) average and $68\%$ C.L. uncertainty, respectively. The solid black lines represent the model defined by Eq. (17). Top: fixed $M_{\text{LIV}}=3\times 10^{17}$ GeV, bottom: fixed $E=10^{19}$ eV. See text for a detailed discussion.
@@ VI Conclusions
We have shown that a subluminal LIV in the photon sector on the mass scale of $M_{\text{LIV}}\sim 10^{16}\,\text{ GeV}$ could be an explanation for the muon puzzle. This parameter range has not yet been tested by data and observations. Moreover, $M_{\text{LIV}}\sim 10^{(15\dots 16)}$ GeV is favored in one of models of Horava-Lifshitz gravity [63]. Additionally, subluminal quartic LIV in the photon sector can appear as a loop effect of LIV in other species [64].
The proposed LIV scenario can be falsified by measuring the ground-level muons spectrum. The latter must match that of a higher energy LI EAS. The obtained numerical value of the LIV mass scale can be tested in a more direct way with the observation of air showers induced by primary photons of energy $\sim 10^{(9\dots 10)}$ GeV. Cosmogenic photons [65] of these energies are produced by ultra-high-energy protons in the Greizen-Zatsepin-Kuzmin process [66, 67]. The corresponding analysis on LIV bounds has been performed in [68]. Thus, a detection of primary photons in this energy range would set a strong constraint $M_{\text{LIV}}>10^{22}$ GeV, which may exclude the LIV explanation of the muon puzzle: this parameter range seems to have almost negligible effect to energy reconstruction.

### v2

@@ I Introduction
One of the most intriguing problems in cosmic-ray physics is that of the “muon excess” in extensive air showers (EAS), or “the muon puzzle”. According to data from a number of experiments, the muon content in EAS exceeds significantly the predictions of models of the development of the shower for the same energy of the primary particle. Solving the muon excess problem is a key to determining accurate models of the development of EAS. Understanding the EAS development is a necessary prerequisite for any reliable study of ultrahigh energy cosmic rays, and at the same time, EAS physics is a tool to study particle interactions in kinematical regimes hard to probe at accelerators.
@@ II.1 An EFT approach to $n=2$ LIV Quantum Electrodynamics
Thus, analyzing LIV for mass scales $10^{(13\dots 16)}$ GeV we can neglect LIV for electrons. In this paper, we neglect LIV for electrons even for higher mass scales since we propose a particular LIV solution of muon puzzle which can be generalized to account for electron LIV parameters.
@@ IV Monte-Carlo simulations
One can see that the MC simulations explicitly justify the use of Eqs. (11)–(14), and (17). The actual particle ratios trend is well captured by our parametric model.
Figure 2: LI–to–LIV ratios for the ground-level average number of particles, test data set (main figures: electrons, inset figures: muons). The gray points and the error bars correspond to MC simulations (proton-induced EASs) average and $68\%$ C.L. uncertainty, respectively. The solid black lines represent the model defined by Eq. (17). Top: fixed $M_{\text{LIV}}=3\times 10^{17}$ GeV, bottom: fixed $E=10^{19}$ eV. See text for a detailed discussion.
@@ VI Conclusions
We have shown that a subluminal LIV in the photon sector on the mass scale of $M_{\text{LIV}}\sim 10^{16}\,\text{ GeV}$ could be an explanation for the muon puzzle. This parameter range has not yet been tested by data and observations. Moreover, $M_{\text{LIV}}\sim 10^{(15\dots 16)}$ GeV is favored in one of models of Horava–Lifshitz gravity [63]. Additionally, subluminal quartic LIV in the photon sector can appear as a loop effect of LIV in other species [64].
The proposed LIV scenario can be falsified by measuring the ground-level muons spectrum. The latter must match that of a higher energy LI EAS. The obtained numerical value of the LIV mass scale can be tested in a more direct way with the observation of air showers induced by primary photons of energy $\sim 10^{(9\dots 10)}$ GeV. Cosmogenic photons [65] of these energies are produced by ultra-high-energy protons in the Greizen–Zatsepin–Kuzmin process [66, 67]. The corresponding analysis on LIV bounds has been performed in [68]. Thus, a detection of primary photons in this energy range would set a strong constraint $M_{\text{LIV}}>10^{22}$ GeV, which may exclude the LIV explanation of the muon puzzle: this parameter range seems to have almost negligible effect to energy reconstruction.